OpenStudy (anonymous):

How many moles of sulfuric acid are needed to produce 57.8 milliliters of water? Water has the density of 0.987 g/mL. Show all steps of your calculation as well as the final answer. NaOH + H2SO4 → Na2SO4 + H2O @nacholover

OpenStudy (anonymous):

57.8ml x 0.987g/ml = 57.05g H2O 57.05g H2O / 18g/mole = 3.17moles H2O 2NaOH + H2SO4 → Na2SO4 + 2H2O.....balanced 3.17moles H2O x (1H2SO4 / 2H2O) = 1.58moles H2SO4

OpenStudy (anonymous):

what do i need to do with NaOH + H2SO4 → Na2SO4 + H2O

OpenStudy (anonymous):

I'm not sure, That was just there with the question, is that the formula?

OpenStudy (anonymous):

who can help me with my question?

OpenStudy (anonymous):

me hold on let me finish his

OpenStudy (anonymous):

Yea I dunno what the formula is

OpenStudy (anonymous):

thats all @Redneck1444 but @sgarciag141 whats your question

OpenStudy (anonymous):

anything else

OpenStudy (anonymous):

Thank you!!! yes one more, I'll post it as a separate question

OpenStudy (anonymous):

Hey im berly in highschool

OpenStudy (anonymous):

well im in th grade so i graduate next year. sorry im late i can help yall anytime.